A​ new, non-evasive, test is to indicate the existance of colorectal cancers in patients who have colorectal cancer with a probability of 0.95. ​However, this test will also show the presense of colorectal cancer in patients who do not have colorectal cancer 22​% of the time.

The chance that a person taking this test has colorectal cancer is 0.09

Find the sensitivity of this test. Use two decimals.

b. Find the specificity of this test. Use two decimals.

c. A person taking this test has tested positive for colorectal cancer. What is the probability​ he/she has colorectal​ cancer? Use four decimals.

d. What is the probability that the person in part c does not have colorectal​ cancer? Use four decimals.
Answer #1

a) sensitivity of this test =P(tested positive|have disease) =0.95

b) specificity of this test =P(tested negative |not have disease) =1-0.22 =0.78

c)

P(tested positive)=P(have disease and tested positive)+P(not have disease and tested positive)
=0.09*0.95+0.91*0.22=0.2857
Therefore P(have disease|tested positive)=0.09*0.95/0.2857=0.2993

d)

P(not have disease given tested positive) =1-P(have disease|tested positive) =1-0.2993 =0.7007